Kathy bicycles 6 km/h faster than Javier. In the same time it takes Javier to bicycle 48 km, Kathy can bicycle 66 km. How fast does each bycyclist travel?
Answer by saifoo.khan(162)

Hello,
k = j+
48/k = 66/k+6

now after solving,
k = 16
So,
kathy = 22
javier = 16
Hope u got your answer.
=============================
If u want free mathlab help, please feel free to contact me.
Good grade gurantee.
Email, einstein.saifoo@gmail.com